OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = (x - 3)/2, what is the value of f[g(-5)]? would this be -7 then

OpenStudy (anonymous):

answer is -5

OpenStudy (anonymous):

f(g(x))=f(g(-5))=2[((-5)-3)/2)]+3

OpenStudy (anonymous):

\[g(-5)=(-5-3)/2=-4\] plug -4 into f(x)=f(-4)\[=-4\times2+3=-5\]

OpenStudy (anonymous):

thx :]